# How to Tell When a Result Will Replicate: Significance and Replication in Distributional Null Hypothesis Tests

### v2

@@ Abstract
There is a well-known problem in Null Hypothesis Significance Testing: many statistically significant results fail to replicate in subsequent experiments. We show that this problem arises because standard ‘point-form null’ significance tests consider only within-experiment but ignore between-experiment variation, and so systematically underestimate the degree of random variation in results. We give an extension to standard significance testing that addresses this problem by analysing both within- and between-experiment variation. This ‘distributional null’ approach does not underestimate experimental variability and so is not overconfident in identifying significance; because this approach addresses between-experiment variation, it gives mathematically coherent estimates for the probability of replication of significant results. Using a large-scale replication dataset (the first ‘Many Labs’ project), we show that many experimental results that appear statistically significant in standard tests are in fact consistent with random variation when both within- and between-experiment variation are taken into account in this approach. Further, grouping experiments in this dataset into ‘predictor-target’ pairs we show that the predicted replication probabilities for target experiments produced in this approach (given predictor experiment results and the sample sizes of the two experiments) are strongly correlated with observed replication rates. Distributional null hypothesis testing thus gives researchers a statistical tool for identifying statistically significant and reliably replicable results. Keywords: Hypothesis Testing; Replication; Significance;
@@
While the logic of evidential significance testing is compelling (a result is surprising when it is extreme, relative to our expectations; a low $p$-value tells us that our result is extreme, relative to the null model) it has become increasingly clear that significance testing, at least as it is typically used, is fundamentally flawed. Perhaps the most important problem with standard significance testing involves replication and the reliability of results, with standard significance testing being overconfident in identifying apparently nonsensical effects, such as telepathy or precognition, as real (parapsychology ‘offers a truly alarming massive case study of how statistics can mislead and be misused’: Diaconis et al.,, 1991), and more generally, with systematic research showing that many statistically significant experimental results in psychology and related fields do not occur reliably in replications (e.g. Camerer et al.,, 2018; Open Science Collaboration et al.,, 2015; Klein et al.,, 2018, 2014), results seen as ‘reflecting an unprecedented level of doubt among practitioners about the reliability of research findings in the field’ (Pashler and Wagenmakers,, 2012). These flaws are also evident in well-known effects of sample size on significance (the observation that the probability of getting a statistically significant result in the standard approach increases with sample size, irrespective of the presence or absence of a true effect; to quote Thompson, (1998): ‘Statistical testing becomes a tautological search for enough participants to achieve statistical significance. If we fail to reject, it is only because we’ve been too lazy to drag in enough participants’) and in the fact that in standard significance tests, null hypotheses are always false (and to quote Cohen, 2016: ‘if the null hypothesis is always false, what’s the big deal about rejecting it?’). Faced with these problems and others (primarily related to hypothesis comparison and optional stopping, e.g. Wagenmakers et al.,, 2018; Wagenmakers,, 2007) there have been increasingly widespread calls for the wholesale abandonment of the significance testing approach (e.g. McShane et al.,, 2019; Amrhein and Greenland,, 2018; Wagenmakers et al.,, 2018; Hunter,, 1997; Carver,, 1978).
@@
Our aim in this paper is to show that these problems all arise from a basic gap in the standard approach to significance testing: the fact that it considers only within-experiment random variation in estimation, while ignoring between-experiment random variation. In the language of measurement analysis in physics, the standard approach considers the precision of the experiment as a measurement tool (the degree of random variation in measurements within this experiment, a factor which depends on sample size) but ignores the accuracy of the experiment as a tool (the degree to which measurements in this experiment agree with those obtained from other experiments measuring the same effect; e.g Mandel,, 2012). This gap means that the standard approach systematically underestimates the degree of random variation or error affecting experimental results, and so is overconfident in identifying results as significant. We address this by describing a model of significance testing that includes both within- and between-experiment random variation; the chance of statistical significance in this ‘distributional’ model does not rise with sample size irrespective of any real effect; this approach does not overestimate the random variation affecting experimental results and so is not overconfident in identifying significance; importantly, this model gives coherent estimates for the probability of replication of significant results in future experiments.
@@
The model we describe mixes statistical inference typically associated with the Frequentist approach with that typically associated with the Bayesian: in section $1$ we give a general sketch and overview of the model, positioning it relative to these two approaches. In section $2$ we derive basic expressions for significance, statistical power and probability of replication in this model in the context of a $t$-test, assuming known between-experiment variation. In section $3$ we give unbiased sample estimators for between-experiment variance of means, give expressions for significance and replication in terms of those sample estimates, and extend these to give simple and practical expressions for significance and replication in terms of the standard $T$ distribution. We also extend this approach to tests of correlation and linear regression, and tests of contingency, and briefly compare this model to a well-known point-form approach to replication (Killeen, 2005a, ; Killeen,, 2007). In section $4$ we test this model against the first ‘Many Labs’ replication dataset (Klein et al.,, 2014), showing that this model’s assumption of between-experiment variation holds, that its judgements of statistical significance are many orders of magnitude more conservative than those obtained in the standard approach, and that its replication probabilities reliably and accurately predict observed rates of successful replication of individual results. In section $5$ we compare this model to Bayesian approaches to hypothesis testing via the JZS Bayesian $t$-test (e.g. Rouder et al.,, 2009), showing that this Bayesian approach gives results almost identical to those of a point-form significance test, at least in the Many Labs dataset. In the final section we address some possible criticisms of this model of significance and replication, and argue that this distributional approach to evidential significance testing will help researchers identify experimental results which are both surprising (statistically significant) and amenable to further investigation (replicable). Rather than abandoning significance testing as a way of identifying surprising and interesting results, researchers should consider using significance testing in its more general distributional form.
1 Overview of the distributional model
@@
The basic distinction between Frequentist and Bayesian approaches to statistical inference lies in the interpretation of probability: the Frequentist sees probability as the limit of relative frequency in sampling, while the Bayesian sees probability as representing degree of belief or uncertainty in estimates. This distinction means that Frequentist statistical inference considers the distribution of measurements in samples, while Bayesian inference considers both this sampling distribution and additional prior distributions (with these priors reflecting degrees of belief or uncertainty in parameters of the sampling distribution). For example: in a simple experiment estimating some unknown parameter $\mu$ from $N$ measurements and assuming Normal error with variance $\sigma^{2}$, the standard Frequentist approach assumes that the sample mean $\overline{X}$ is distributed as
@@
and, assuming for simplicity that the sample variance $S^{2}$ is an accurate estimate of $\sigma^{2}$, a given result is statistically significant relative to the null hypothesis $H_{0}:\mu=0$ at level $\alpha$ when
$P(\overline{X}|H_{0})=2\Phi\left(-\frac{|\overline{X}|}{S/\sqrt{N}}\right)<\alpha$
where $P(\overline{X}|H_{0})$ represents the probability of observing a result equal to or more extreme than $\overline{X}$ given $H_{0}$, and $\Phi$ represents the cumulative Normal distribution.
A Bayesian approach in this situation, by contrast, assumes a prior distribution of $\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ with $\mu_{0}$ representing an estimate for the parameter of interest and $\sigma_{0}^{2}$ representing uncertainty or degree of belief in that estimate, so that
@@
and so taking $H_{0}:\mu_{0}=0$ (and assuming again that $S^{2}$ accurately estimates $\sigma^{2}$) this gives
$P(\overline{X}|H_{0})=2\Phi\left(-\frac{|\overline{X}|}{\sqrt{S^{2}/N+\sigma_{0}^{2}}}\right)$
In fact, however, Bayesians are typically more interested in the updated or posterior distribution
@@
$\overline{X}\sim\mathcal{N}(\mu_{0},S^{2}/N+\sigma_{0}^{2})$
just as in the Bayesian approach. Given this distribution, and estimating $\sigma_{0}^{2}$ by the sample variance or heterogeneity of means across similar experiments, $S^{2}_{0}$ (just as $\sigma^{2}$ is estimated by the sample variance of measurements within a single experiment, $S^{2}$) we can assess statistical significance relative to both within- and between-experiment variance: a given result will be statistically significant relative to the null hypothesis $H_{0}:\mu_{0}=0$ when
@@
We can also produce a posterior distribution $\mu\sim\mathcal{N}(m|\mu_{N},\sigma_{N}^{2})$ by applying Bayes’ theorem as before, giving a distribution for hypotheses given data of $P(H|\overline{X})$. In this distributional approach, both directions of inference (probability of data given hypothesis $P(\overline{X}|H)$ and probability of hypotheses given data $P(H|\overline{X})$) can be applied within the Frequentist perspective, with the difference between Frequentist and Bayesian viewpoints arising in the interpretation of $\sigma_{0}^{2}$ (the Frequentist interpreting it as representing between-experiment variance of experimental means and the Bayesian as representing degree of prior belief or uncertainty about the population mean). Given the same experimental result $\overline{X}$ and the same values for $\mu_{0}$ and $\sigma_{0}^{2}$, however interpreted, in this approach Frequentists and Bayesians will make the same judgement of statistical significance and will infer the same updated distribution for $\mu$.
Our position, then, is that Bayesian and Frequentist approaches to statistical inference differ in their interpretation of probability, but not in the analyses that can be applied to mathematical representations of probability in terms of statistical distributions. It happens that analyses which estimate $P(\overline{X}|H)$ are primarily associated with the Frequentist viewpoint (most likely because estimates $P(H|\overline{X})$ are uninformative under a point-form null hypothesis: when the probability density of $H$ is massed at a single point, $P(H|\overline{X})=P(H)$ necessarily holds) while those which estimate $P(H|\overline{X})$ are primarily associated with the Bayesian approach.111 Although some Bayesian proposals do estimate $P(\overline{X}|H)$, typically as part of a prior check of validity before Bayesian inference (e.g. Dass and Berger,, 2003; Gabry et al.,, 2019; van de Schoot et al.,, 2021; Schad et al.,, 2021). Our distributional model, by contrast, makes use of both $P(\overline{X}|H)$ and $P(H|\overline{X})$ throughout analysis, to estimate both the significance of new results and the probability of obtaining similar results in future replications.
Some readers may initially see this model as a confusing or contradictory mix of Bayesian and Frequentist interpretations of probability, a perception arising from the fact that the model involves both inference from hypothesis to data (typically Frequentist) and inference from data to hypothesis (typically Bayesian). This perception is incorrect, because these two forms of conditional inference apply to mathematical representations of probability as statistical distributions, not to interpretations of probability (as limiting relative frequencies or as degrees of belief). As such these two forms of inference can be applied simultaneously, with no contradiction, under any interpretation where probabilities are represented as distributions. We expand this initial sketch to give a full and detailed presentation of our distributional model below.

### v3

@@ Abstract
There is a well-known problem in Null Hypothesis Significance Testing: many statistically significant results fail to replicate in subsequent experiments. We show that this problem arises because standard ‘point-form null’ significance tests consider only within-experiment but ignore between-experiment variation, and so systematically underestimate the degree of random variation in results. We give an extension to standard significance testing that addresses this problem by analysing both within- and between-experiment variation. This ‘distributional null’ approach does not underestimate experimental variability and so is not overconfident in identifying significance; because this approach addresses between-experiment variation, it gives mathematically coherent estimates for the probability of replication of significant results. Using a large-scale replication dataset (the first ‘Many Labs’ project), we show that many experimental results that appear statistically significant in standard tests are in fact consistent with random variation when both within- and between-experiment variation are taken into account in this approach. Further, grouping experiments in this dataset into ‘predictor-target’ pairs we show that the predicted replication probabilities for target experiments produced in this approach (given predictor experiment results and the sample sizes of the two experiments) are strongly correlated with observed replication rates. Distributional null hypothesis testing thus gives researchers a statistical tool for identifying statistically significant and reliably replicable results.
Keywords: Hypothesis Testing; Replication; Significance
@@
While the logic of evidential significance testing is compelling (a result is surprising when it is extreme, relative to our expectations; a low $p$-value tells us that our result is extreme, relative to the null model) it has become increasingly clear that significance testing, at least as it is typically used, is fundamentally flawed. Perhaps the most important problem with standard significance testing involves replication and the reliability of results, with standard significance testing being overconfident in identifying nonsensical effects, such as telepathy or precognition, as real (parapsychology ‘offers a truly alarming massive case study of how statistics can mislead and be misused’: Diaconis et al.,, 1991), and more generally, with systematic research showing that many statistically significant experimental results in psychology and related fields do not occur reliably in replications (e.g. Camerer et al.,, 2018; Open Science Collaboration et al.,, 2015; Klein et al.,, 2018, 2014), results seen as ‘reflecting an unprecedented level of doubt among practitioners about the reliability of research findings in the field’ (Pashler and Wagenmakers,, 2012). These flaws are also evident in well-known effects of sample size on significance (the observation that the probability of getting a statistically significant result in the standard approach increases with sample size, irrespective of the presence or absence of a true effect; to quote Thompson, (1998): ‘Statistical testing becomes a tautological search for enough participants to achieve statistical significance. If we fail to reject, it is only because we’ve been too lazy to drag in enough participants’) and in the fact that in standard significance tests, null hypotheses are always false (and to quote Cohen, 2016: ‘if the null hypothesis is always false, what’s the big deal about rejecting it?’). Faced with these problems and others (primarily related to hypothesis comparison and optional stopping, e.g. Wagenmakers et al.,, 2018; Wagenmakers,, 2007) there have been increasingly widespread calls for the wholesale abandonment of the significance testing approach (e.g. McShane et al.,, 2019; Amrhein and Greenland,, 2018; Wagenmakers et al.,, 2018; Hunter,, 1997; Carver,, 1978).
@@
We see these problems as arising from a basic gap in the standard approach to significance testing: the fact that it considers only within-experiment random variation in estimation, while ignoring between-experiment random variation. In the language of measurement analysis in physics, the standard approach considers the precision of the experiment as a measurement tool (the degree of random variation in measurements within this experiment, a factor which depends on sample size) but ignores the accuracy of the experiment as a tool (the degree to which measurements in this experiment agree with those obtained from other experiments measuring the same effect; e.g Mandel,, 2012). This gap means that the standard approach systematically underestimates the degree of random variation or error affecting experimental results, and so is overconfident in identifying results as significant. We address this by describing a model of significance testing that includes both within- and between-experiment random variation; the chance of statistical significance in this ‘distributional’ model does not rise with sample size irrespective of any real effect; this approach does not overestimate the random variation affecting experimental results and so is not overconfident in identifying significance; importantly, this model gives coherent estimates for the probability of replication of significant results in future experiments.
The idea of between-experiment variation is, of course, well known, and is referred to as ‘heterogeneity’ in the meta-analytic and systematic review literatures. Researchers carrying out a systematic review of a set of studies on a given result will typically distinguish between clinical heterogeneity or diversity (between-experiment variation in the participants, interventions and outcomes studied), methodological heterogeneity (variation in study design, measurement tools, and control of confounding factors), and outcome heterogeneity (variation in the precise effect being evaluated in each experiment) with outcome heterogeneity arising as a consequence of clinical or methodological differences between studies (see e.g. Chandler et al.,, 2019). One strand of research in this area is causal, and focuses on ‘hidden moderators’: particular differences in study design, participant population, or measured outcome that are hypothesised to produce the observed difference in experimental results (Kenny and Judd,, 2019; Tackett et al.,, 2017). Such hidden moderators are particularly plausible when considering the heterogeneity of psychological results across cultures: studies designed in a particular cultural setting may implicitly depend on some specific aspect of that culture, with participants from a different cultural background giving obviously different responses. A clear example of this arises in our analysis of results from the Many Labs $1$ project (Klein et al.,, 2014), which ran studies replicating $16$ classic cognitive and social psychology results at $36$ different sites across the world (see below). One of these studies examined the effect of ‘anchoring’ on responses to numerical questions by asking participants to estimate the height of Mount Everest: where estimates given by participants at US sites were typically quite far from the correct height, participants at some international sites gave estimates that were only a few metres from the correct height (indicating a significant cross-cultural difference in participant knowledge about the actual height of Mount Everest).
In the ‘hidden moderator’ approach, heterogeneity in experimental results arises as a consequence of one or two major and identifiable differences in experimental method or participant population: in this perspective heterogeneity can be controlled by careful design. A contrasting ‘random-effects’ approach sees heterogeneity as arising, not from one or two major differences between experiments, but instead from the combined effect of many small and random differences: from the Central Limit theorem, this combined effect will produce Normally distributed variability between experimental means, just as the combined effect of small and random differences between individual participants produces Normally distributed variability in participant responses. This assumption of Normality allows us to use standard statistical methods to estimate between-experiment variance (see e.g. Hedges,, 1983; Viechtbauer,, 2005; Veroniki et al.,, 2016): in systematic reviews this type of random-effects analysis is used to increase precision in our overall estimate of the size of an effect of interest by combining results from different studies, giving more weight to results with lower within-experiment variance. [^1] Note that in presenting our approach we use the within-experiment/between-experiment variance terminology, rather than distinguishing between ‘variance’ and ‘heterogeneity’. We make this terminological choice because, following this random-effects approach, we see ‘heterogeneity’ as just another form of statistical variance, and to stress that between-experiment variance arises even when experiments are not obviously heterogeneous (do not have obvious differences in participant population or methodology).
Rather than using between-experiment variance to increase the precision of our total estimate for an effect (parameter estimation via post-hoc meta-analysis of existing results, as in systematic review), our approach uses between-experiment variance to construct an a priori null hypothesis designed to include both within- and between-experiment variance when assessing the statistical significance of new experimental results. This distributional null hypothesis assumes, not that the true effect is 00 across all experiments (as in standard or ‘point-form’ null hypothesis tests), but instead that the true effect is randomly distributed across experiments around a mean of 00 (with a given level of between-experiment variance). A statistically significant experimental result, in this distributional approach, is one which is inconsistent with an hypothesised mean of 00 even under this assumption of between-experiment variance. We can illustrate this point using a recent ‘W-Boson mass anomaly’ in particle physics, arising from 2022 W-Boson mass measurements by the Collider Detector at Fermilab collaboration (CDF Collaboration et al.,, 2022). These measurements were significantly higher than the Standard Model prediction for that particle mass and were inconsistent with previous experimental results at a $p<10^{-6}$ significance level in standard statistical tests. This anomaly was seen as suggesting new physics beyond the Standard Model, such as new particles or interactions that modify the $W$-boson’s mass (e.g. Endo and Mishima,, 2022; Ghorbani and Ghorbani,, 2022; Du et al.,, 2023; Zhang and Feng,, 2023). Using our model of significance to analyse estimates of the $W$-Boson mass across a range of particle physics experiments (Costello and Watts,, 2023) we found that these estimates were distributed with Normal between-experiment error: the anomalous CDF measurements were no longer statistically significant when the effect of this between-experiment random variation was taken into account.
@@
The model we describe mixes a form of statistical inference typically associated with the Frequentist approach (significance testing) with a form typically associated with the Bayesian approach (prior distributions over effect sizes): in section $1$ we give a general sketch and overview of the model, positioning it relative to these two approaches. In section $2$ we derive basic expressions for significance, statistical power and probability of replication in this model in the context of a $t$ test, assuming known between-experiment variation. In section $3$ we give unbiased sample estimators for between-experiment variance of means (following a standard meta-analytic approach), give expressions for significance and replication in terms of those sample estimates, and extend these to give simple and practical expressions for significance and replication in terms of the standard $t$ distribution. We also extend this approach to tests of correlation and linear regression, and tests of contingency, and briefly compare this model to a well-known point-form approach to replication (Killeen, 2005a, ; Killeen,, 2007). In section $4$ we test this model against the first ‘Many Labs’ replication dataset (Klein et al.,, 2014), showing that this model’s assumption of between-experiment variation holds, that its judgements of statistical significance are many orders of magnitude more conservative than those obtained in the standard approach, and that its replication probabilities reliably and accurately predict observed rates of successful replication of individual results. In section $5$ we compare this model to Bayesian approaches to hypothesis testing via the JZS Bayesian $t$ test (e.g. Rouder et al.,, 2009), showing that this Bayesian approach gives results almost identical to those of a point-form significance test, at least in the Many Labs dataset. In the final section we address some possible criticisms of this model of significance and replication, and argue that this distributional approach to evidential significance testing will help researchers identify experimental results which are both surprising (statistically significant) and amenable to further investigation (replicable). Rather than abandoning significance testing as a way of identifying surprising and interesting results, researchers should consider using significance testing in a more general distributional form.
1 Sketch of the distributional model
We begin by sketching our model in very broad strokes, by placing it relative to standard Frequentist and Bayesian approaches to statistics. This preliminary sketch necessarily ignores many important details: these are addressed in the full model presentation, below.
@@ 1 Sketch of the distributional model
The basic distinction between Frequentist and Bayesian approaches to statistical inference lies in the interpretation of probability: the Frequentist sees probability as the limit of relative frequency in sampling, while the Bayesian sees probability as representing degree of belief or uncertainty in estimates. This distinction means that Frequentist statistical inference considers the distribution of measurements in samples, while Bayesian inference considers both this sampling distribution and additional prior distributions (with these priors reflecting degrees of belief or uncertainty in parameters of the sampling distribution). For example: in a simple experiment estimating some unknown parameter $\mu$ from $N$ measurements and assuming Normal error with unknown variance $\sigma^{2}$, the standard Frequentist approach assumes that the sample mean $\overline{X}$ is distributed as
@@
(the Normal error model). Defining the test statistic
$t=\frac{|\overline{X}|}{S/\sqrt{N}}$
and taking the sample standard deviation $S$ to be an accurate estimate of the model standard deviation $\sigma$, then a given result will be statistically significant relative to the null hypothesis $H_{0}:\mu=0$ at level $\alpha$ in a two-sided test when
$P(\overline{X}|H_{0})=2\Phi\left(\frac{-|\overline{X}|}{S/\sqrt{N}}\right)<\alpha$
where $P(\overline{X}|H_{0})$ represents the probability of observing a result equal to or more extreme than $\overline{X}$ given $H_{0}$ and $\Phi$ represents the cumulative Normal distribution. This is the $Z$ test for significance, and is valid only when we can reasonably assume that $S$ accurately estimates $\sigma$ (an assumption which only holds for large $N$). We can generalise this approach by considering the distribution of the test statistic $t$ under the Normal error model: for an experiment with $\nu$ degrees of freedom, this statistic has the cumulative $t$ distribution $T_{\nu}$, and a result will be statistically significant relative $H_{0}:\mu=0$ at level $\alpha$ in a two-sided test when
$2T_{\nu}\left(\frac{-|\overline{X}|}{S/\sqrt{N}}\right)<\alpha$
This distribution $T_{\nu}$ approaches $\Phi$ as the sample size $N$ increases (that is, as $S$ more accurately estimates $\sigma$).
Where this ‘Frequentist’ approach assumes a null hypothesis with $\mu=0$, the Bayesian approach assumes a prior distribution $\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ with $\mu_{0}$ representing a prior estimate for the parameter of interest and $\sigma_{0}^{2}$ representing prior uncertainty or degree of belief in that estimate. Here our statistical model is
@@
and so taking $H_{0}:\mu_{0}=0$, if we assume that $S^{2}$ accurately estimates $\sigma^{2}$ (as before, a reasonable assumption for large $N$) then we have
$P(\overline{X}|H_{0})=2\Phi\left(\frac{-|\overline{X}|}{\sqrt{S^{2}/N+\sigma_{0}^{2}}}\right)$
Bayesians, however, are typically more interested in the updated or posterior distribution
@@
$\overline{X}\sim\mathcal{N}(\mu_{0},\sigma^{2}/N+\sigma_{0}^{2})$
just as in the Bayesian approach. If we assume that $\sigma$ is accurately estimated by the sample variance $S$ and that $\sigma_{0}^{2}$ by is accurately estimated by the sample variance or heterogeneity of means across similar experiments, $S^{2}_{0}$, we can assess statistical significance relative to both within- and between-experiment variance: a given result will be statistically significant relative to the null hypothesis $H_{0}:\mu_{0}=0$ when
@@
Just as before, we can also consider the distribution of the test statistic here under our statistical model for $\overline{X}$: it turns out that a particular transformation of this statistic follows the $t$ distribution with $\nu$ degrees of freedom as before, and so we can also test for significance against $T_{\nu}$ (see the next section).
We can also produce a posterior distribution $\mu\sim\mathcal{N}(m|\mu_{N},\sigma_{N}^{2})$ by applying Bayes’ theorem as before, giving a distribution for hypotheses given data of $P(H|\overline{X})$. In this distributional approach, both directions of inference (‘Frequentist’ estimation of the probability of data given hypothesis $P(\overline{X}|H)$, and ‘Bayesian’ estimation of the probability of hypotheses given data $P(H|\overline{X})$) can be applied simultaneously, with the difference between Frequentist and Bayesian viewpoints arising in the interpretation of $\sigma_{0}^{2}$ (the Frequentist interpreting it as representing between-experiment variance of experimental means and the Bayesian as representing degree of prior belief or uncertainty about the population mean). Given the same experimental result $\overline{X}$ and the same values for $\mu_{0}$ and $\sigma_{0}^{2}$, however interpreted, in this approach Frequentists and Bayesians will make the same judgement of statistical significance and will infer the same updated distribution for $\mu$.
Our position, then, is that Bayesian and Frequentist approaches to statistical inference differ in their interpretation of probability, but not in the analyses that can be applied to mathematical representations of probability in terms of statistical distributions. It happens that analyses which estimate $P(\overline{X}|H)$ are primarily associated with the Frequentist viewpoint (most likely because estimates $P(H|\overline{X})$ are uninformative under a point-form null hypothesis: when the probability density of $H$ is massed at a single point, $P(H|\overline{X})=P(H)$ necessarily holds) while those which estimate $P(H|\overline{X})$ are primarily associated with the Bayesian approach.222 Although some Bayesian proposals do estimate $P(\overline{X}|H)$, typically as part of a prior check of validity before Bayesian inference (e.g. Dass and Berger,, 2003; Gabry et al.,, 2019; van de Schoot et al.,, 2021; Schad et al.,, 2021). Our distributional model, by contrast, makes use of both $P(\overline{X}|H)$ and $P(H|\overline{X})$ throughout analysis, to estimate both the significance of new results and the probability of obtaining similar results in future replications.